# Variational Improvement of the Hartree-Fock Approach to the 2D Hubbard Model

### v1

@@ Abstract
We consider a refinement of the usual Hartree-Fock method applied the 2D Hubbard model, in Nambu spinor formulation. The new element is the addition of a “condensate inducing” term proportional to a variational parameter $h$ to the Hartree-Fock Hamiltonian, which generates an s- or d-wave condensate. This modified Hartree-Fock Hamiltonian is used only to generate variational trial states; energy expectation values are computed in the full two-dimensional Hubbard Hamiltonian with no modification. It is found that there exist trial states with non-vanishing condensates that are significantly lower in energy than the standard Hartree-Fock ground states, in some range of the $U/t$-density plane.
@@
There is still some question about whether, and for which parameter values, the ground state of the two-dimensional Hubbard model exhibits the d-wave condensation observed in cuprate superconductors. The model is difficult to solve convincingly due to the sign problem, and all existing methods seem to have their limitations, both theoretical and practical. The earliest approach, dating to 1966, is the Hartree-Fock method applied to the Hubbard model, and this venerable approach has a very large literature, a sample of which is Penn ; *Poilblanc; *Zaanen; *Machida; *Schulz1; *Schulz2; *Ichimura; *Verges; *Inui; *Dasgupta; *Xu. Two contributions of ours along these lines are in Matsuyama:2022kam ; Matsuyama:2022dtd . The strengths and limitations of the Hartree-Fock method are well known, and are described in a number of reviews Scalettar ; Powell ; Lechermann ; Imada ; Fazekas . Like any mean field approach the method neglects correlations, and in addition uses a single Slater determinant to approximate what is surely a more complex ground state. Nevertheless, this method was successful in predicting stripe patterns (Zaanen and Gunnarsson, Poilblanc and Rice in Zaanen ) later observed in experiment Emery , as well as the emergence of ferromagnetism and the relation to the Stoner criterion. So it is of interest to see how far one can go with a single Slater determinant, in particular whether the existing Hartree-Fock approach can be significantly improved, and whether an improved approach can tell us anything about d-wave condensation and superconductivity.
@@
We begin with the 2D Hubbard model, in Nambu spinor formalism Matsuyama:2022dtd
@@
$\psi(x)$ $=$ $\left[\begin{array}[]{c}c_{\uparrow}(x)\cr c^{\dagger}_{\downarrow}(x)\end{array}\right]~{}~{}~{},~{}~{}~{}\psi^{\dagger}(x)=[c^{\dagger}_{\uparrow}(x),c_{\downarrow}(x)]\ ,$ (4)
@@
Note that the Fock vacuum $|0\rangle$ in Nambu formulation, annihilated by all $\psi$ operators, differs from the no-electron Fock vacuum in standard formulation. $|\Omega\rangle$ is not an eigenstate of particle number, in the Nambu formulation, but it is an eigenstate of the difference of electron spin up/down number operators $N_{\uparrow}-N_{\downarrow}$. Setting $M=L^{2}$ restricts to states with an absence of overall magnetic moment (equal numbers of spin up and down), with the expectation value of electron density controlled by the chemical potential $\mu$. In the Hartree-Fock approach the one-particle wave functions $\phi_{i}(x_{i},\alpha_{i})$ are to be determined self-consistently, and defining
$\rho(x,\alpha\beta)$ $=$ $\langle\Omega|\psi^{\dagger}_{\alpha}(x)\psi_{\beta}(x)|\Omega\rangle\ ,$ (7)
@@
$H_{HF}(\Omega)$ (8)
$=$ $-t\sum_{<xy>}(\psi_{1}^{\dagger}(x)\psi_{1}(y)-\psi_{2}^{\dagger}(x)\psi_{2}(y))$
$-\mu\sum_{x}(\psi_{1}^{\dagger}(x)\psi_{1}(x)-\psi_{2}^{\dagger}(x)\psi_{2}(x))$
$+U\sum_{x}\bigg{\{}[1-\rho(x,22)]\psi^{\dagger}_{1}(x)\psi_{1}(x)-\rho(x,11)\psi^{\dagger}_{2}(x)\psi_{2}(x)$
$\qquad\qquad+\rho(x,12)\psi^{\dagger}_{2}\psi_{1}+\rho(x,21)\psi_{1}^{\dagger}\psi_{2}\bigg{\}}$
@@
There is no obvious evidence of a condensate in the ground state of the Hartree-Fock Hamiltonian, in either conventional Penn ; *Poilblanc; *Zaanen; *Machida; *Schulz1; *Schulz2; *Ichimura; *Verges; *Inui; *Dasgupta; *Xu or Nambu spinor Matsuyama:2022dtd formulation. In general, in Nambu formulation, a condensate exists if there are one-particle states in the Slater determinant such that $\phi_{i}(x,\alpha)\neq 0$ for both indices $\alpha=1,2$ (see eq. (13) below). But it is clear from inspection of (6), (7), (8) that Hartree-Fock self-consistency allows that each $\phi_{i}(x,\alpha)$ is nonzero for only one of the two indices $\alpha=1,2$ (and $\rho(x,12)=\rho^{*}(x,21)=0$). In that case there is no condensate in this mean field approach.
@@
where $\hat{e}_{x},\hat{e}_{y}$ are unit vectors in the $x,y$ directions respectively, $q=\pm 1$, and $h$ is a variational parameter. We will denote the modified Hartree-Fock Hamiltonian as $H_{HFq}$. If we make the special assumption that the $\rho(x,\alpha\beta)=\rho(\alpha\beta)$ are position independent (an assumption which is simply not true in the stripe and antiferromagnetic phases) then it is not hard to see that the ground state of $H_{HF-}$ will have a d-wave condensate in the neighborhood of the Fermi surface, even in the limit $h\rightarrow 0$. This result, which follows from degenerate perturbation theory, can be seen as follows: Translation invariance of $\rho$ implies that $[H]$ is diagonalized by momentum eigenstates, and on an $L\times L$ lattice, in the basis
@@
the matrix $[H]$ is decomposed, at each $k$, into blocks
$[H]_{k}=\left[\begin{array}[]{cc}\epsilon^{1}_{k}&2h(\cos(k_{x})-\cos(k_{y}))\cr 2h(\cos(k_{x})-\cos(k_{y}))&\epsilon^{2}_{k}\end{array}\right]\ ,$ (11)
@@
The pairing condensate in momentum space is
$=$ ${1\over L^{2}}\sum_{i=1}^{L^{2}}\phi_{i}^{*}(p,1)\phi_{i}(p,2)\ .$
As $h\rightarrow 0$ the wavefunctions (10) are eigenstates of $[H]_{k}$, in which case $P(\mathbf{p})=0$, and there is no condensate except at momenta where $\epsilon^{1}_{k}=\epsilon^{2}_{k}$. For $\langle N_{\uparrow}-N_{\downarrow}\rangle=0$, and denoting the electron number density on an $L\times L$ periodic lattice as
$f$ $=$ $\langle N_{\uparrow}+N_{\downarrow}\rangle/L^{2}$ (14)
$=$ $\sum_{x}\langle\Omega|\psi^{\dagger}_{1}(x)\psi_{1}(x)-\psi^{\dagger}_{2}(x)\psi_{2}(x)+1|\Omega\rangle$
$=$ ${1\over L^{2}}\sum_{x}\{\rho(x,11)-\rho(x,22)+1\}\ ,$
the condition is
$2t((\cos(k_{x})+\cos(k_{y}))-\frac{1}{2}Uf+\mu=0\ ,$ (15)
which we refer to as the Fermi surface. Precisely on the Fermi surface, degenerate perturbation theory gives us new eigenstates $\phi^{\prime}_{\pm}(k)=(1/\sqrt{2})(\phi_{1}(k)\pm\phi_{2}(k))$, resulting in a non-zero condensate. This is sufficient to change the condensate pattern dramatically for any finite $h$, no matter how small, and as $h\rightarrow 0$ we find Matsuyama:2022dtd
@@
$P(\mathbf{k})=\left\{\begin{array}[]{cl}\frac{1}{2}\text{sign}(\cos k_{x}-\cos k_{y})&\mathbf{k}\in\text{Fermi surface}\cr 0&\text{otherwise}\end{array}\right.\ ,$ (18)
@@
a result which is strongly reminiscent of d-wave pairing. Since this seems to be a case where the slightest perturbation breaks a ground state degeneracy, even as the strength of the breaking terms goes to zero, it is tempting to regard this situation as an example of spontaneous symmetry breaking. In fact this argument was already presented in Matsuyama:2022dtd , but there are two objections. First, the densities $\rho(x,\alpha\beta)$ are not necessarily translation invariant, e.g. in the stripe phase, yet there can be superconductivity in that phase. But the more serious objection is that if we take the $h\rightarrow 0$ limit, then by this argument the condensate only exists exactly at the Fermi surface. In that case the condensate density vanishes in the infinite volume limit.
These objections suggest a change in philosophy: why should we take the $h\rightarrow 0$ limit? The real objective is to find the best possible approximation, within the limitations dictated by a single Slater determinant, to the ground state of the full $h$-independent Hubbard Hamiltonian of (1). Then the question is whether mixing one-particle states in the neighborhood of the Fermi surface, which is what the condensate-inducing term achieves, would lower the expectation value of the Hubbard Hamiltonian. If so, then since the mixed states would have non-zero amplitudes for both $\alpha$ index values, a condensate is obtained. So let us regard the ground state of $H_{HFq}$, denoted $|\Omega_{h}\rangle$, as a trial wave functional, dependent on the variational parameter $h$. If energy expectation value $\mathcal{E}(h)=\langle\Omega_{h}|H|\Omega_{h}\rangle$ of the full Hubbard Hamiltonian is minimized at some $h>0$, then by the rules of the variational approach the ground state $|\Omega_{h}\rangle$ is a better approximation, as compared to the Hartree-Fock ground state $|\Omega_{0}\rangle$, to the true ground state of the Hubbard Hamiltonian. Then the procedure is to choose some set of parameters $\{U,\mu\}$ (setting $t=1$ for numerical work), and at each $U,\mu$ and variational parameter $h$ we compare $\mathcal{E}(0)$ to (i) $\mathcal{E}(h)$ for the ground state $|\Omega_{h}\rangle$ of $H_{HF-}$, associated with a d-wave condensate; and (ii) $\mathcal{E}(h)$ for the ground state $|\Omega_{h}\rangle$ of $H_{HF+}$, associated with an s-wave condensate. If $h>0$ is energetically favorable, in cases (i) and/or (ii), we choose the value of $h$ and sign of $q$ which gives the lowest $\mathcal{E}$. Whichever ground state has the lowest $\mathcal{E}$ is the state which is favored at the given $U,\mu$, at the electron density $f$ in (14). We again emphasize that the $h$ parameter is absent in the Hubbard Hamiltonian $H$, and only appears as a label of the choice of trial ground state.

### v3

# A Variational Improvement of the Hartree-Fock Approach to the 2D Hubbard Model

@@ Abstract
We consider a refinement of the usual Hartree-Fock method applied to the 2D Hubbard model, in Nambu spinor formulation. The new element is the addition of a “condensate inducing” term proportional to a variational parameter $h$ to the Hartree-Fock Hamiltonian, which generates an s- or d-wave condensate at zero temperature. This modified Hartree-Fock Hamiltonian is used only to generate variational trial states; energy expectation values are computed in the full two-dimensional Hubbard Hamiltonian with no modification. It is found that there exist trial states with non-vanishing condensates which are lower in energy than the standard Hartree-Fock ground states. However, these lower energy condensate states exist only in a spatially inhomogeneous (stripe) phase. No lowering of energy, relative to the Hartree-Fock ground state, is found in the spatially homogenous region of the $U-$density phase plane.
I Introduction
@@
There is still some question about whether, and for which parameter values, the ground state of the two-dimensional Hubbard model exhibits the d-wave condensation observed in cuprate superconductors. The model is difficult to solve convincingly due to the sign problem, and all existing methods seem to have their limitations, both theoretical and practical. The earliest approach, dating to 1966, is the Hartree-Fock method applied to the Hubbard model, and this venerable approach has a very large literature, a sample of which is Penn ; *Hirsch; *Poilblanc; *Zaanen; *Machida; *Schulz1; *Schulz2; *Ichimura; *Verges; *Inui; *Dasgupta; *Xu. Two contributions of ours along these lines are in Matsuyama:2022kam ; Matsuyama:2022dtd . The strengths and limitations of the Hartree-Fock method are well known, and are described in a number of reviews Scalettar ; Powell ; Lechermann ; Imada ; Fazekas . Like any mean field approach the method neglects correlations, and in addition uses a single Slater determinant to approximate what is surely a more complex ground state. Nevertheless, this method was successful in predicting stripe patterns (Zaanen and Gunnarsson, Poilblanc and Rice in Zaanen ) later observed in experiment Emery , as well as the emergence of ferromagnetism and the relation to the Stoner criterion. So it is of interest to see how far one can go with a single Slater determinant, in particular whether the existing Hartree-Fock approach can be significantly improved, and whether an improved approach can tell us anything about d-wave condensation and superconductivity.
In this article we consider going a little beyond the standard Hartree-Fock approach by incorporating a variational element. The idea is to modify the Hartree-Fock Hamiltonian by adding a term proportional to a variational parameter $h$ which induces a d- or s-wave condensate, and then regarding the ground state of the modified Hartree-Fock Hamiltonian as a trial state for the unmodified Hubbard Hamiltonian. The energy expectation values of the unmodified Hubbard Hamiltonian, at each $h$, are evaluated in the trial states, and it will be seen that in some regions of the $U$-density plane the trial states, with a non-vanishing condensate, are lower in energy than the standard Hartree-Fock ground state. It is found that spatial inhomogeneity in spin expectation values (“stripes” and other patterns) is correlated with the lowering of energy by d- or s-wave condensates. No condensation is found in the spatially homogeneous phase, at least in this variational approach.
II Hartree-Fock in Nambu spinor formalism
@@
We begin with a review of some formalities. The 2D Hubbard model in Nambu spinor formalism is Matsuyama:2022dtd ,
@@
$\psi(x)$ $=$ $\left[\begin{array}[]{c}c_{\uparrow}(x)\cr\cr c^{\dagger}_{\downarrow}(x)\end{array}\right]~{}~{}~{},~{}~{}~{}\psi^{\dagger}(x)=[c^{\dagger}_{\uparrow}(x),c_{\downarrow}(x)]\ ,$ (5)
@@
Note that the Fock vacuum $|0\rangle$ in Nambu formulation, annihilated by all $\psi$ operators, differs from the no-electron Fock vacuum in standard formulation. $|\Omega\rangle$ is not an eigenstate of particle number, in the Nambu formulation, but it is an eigenstate of the difference of electron spin up/down number operators $N_{\uparrow}-N_{\downarrow}$. Setting $M=L^{2}$ in (7) restricts to states with an absence of overall magnetic moment (equal numbers of spin up and down), with the expectation value of electron density controlled by the chemical potential $\mu$. In the Hartree-Fock approach the one-particle wave functions $\phi_{i}(x_{i},\alpha_{i})$ are to be determined self-consistently, and defining
$\rho(x,\alpha\beta)$ $=$ $\langle\Omega|\psi^{\dagger}_{\alpha}(x)\psi_{\beta}(x)|\Omega\rangle$ (8)
$=$ $\sum_{i}\phi^{*}_{i}(x,\alpha)\phi_{i}(x,\beta)\ ,$
@@
$H_{HF}(\Omega)$ $=$ $-t\sum_{<xy>}(\psi_{1}^{\dagger}(x)\psi_{1}(y)-\psi_{2}^{\dagger}(x)\psi_{2}(y))-\mu\sum_{x}(\psi_{1}^{\dagger}(x)\psi_{1}(x)+1-\psi_{2}^{\dagger}(x)\psi_{2}(x))$ (9)
$+U\sum_{x}\bigg{\{}[1-\rho(x,22)]\psi^{\dagger}_{1}(x)\psi_{1}(x)-\rho(x,11)\psi^{\dagger}_{2}(x)\psi_{2}(x)+\rho(x,12)\psi^{\dagger}_{2}\psi_{1}+\rho(x,21)\psi_{1}^{\dagger}\psi_{2}\bigg{\}}$
$+U\sum_{x}\{\rho(x,11)\rho(x,22)-\rho(x,12)\rho(x,21)\}$
@@
Then the one-particle wavefunctions in the Slater determinant satisfy
$[H]_{x\alpha,y\beta}\phi_{i}(y,\beta)=E_{i}\phi_{i}(x,\alpha)\ .$ (10)
When the one particle energies $E_{i}$ are arranged, for convenience, in order of increasing energy, then the Slater determinant state (7) with $M=L^{2}$ is the zero temperature ground state of the Hubbard model in the Hartree-Fock approximation. The pairing condensate in momentum space is
$=$ ${1\over L^{2}}\sum_{i=1}^{L^{2}}\phi_{i}^{*}(p,1)\phi_{i}(p,2)$
on an $L\times L$ lattice volume.
There is no obvious evidence of a condensate in the ground state of the Hartree-Fock Hamiltonian, in either conventional Penn ; *Poilblanc; *Zaanen; *Machida; *Schulz1; *Schulz2; *Ichimura; *Verges; *Inui; *Dasgupta; *Xu or Nambu spinor Matsuyama:2022dtd formulation. In general, in Nambu formulation, a condensate exists if there are one-particle states in the Slater determinant such that $\phi_{i}(x,\alpha)\neq 0$ for both indices $\alpha=1,2$, as we see in eq. (11). But it is clear from inspection of (7), (8), (9) that Hartree-Fock self-consistency allows that each $\phi_{i}(x,\alpha)$ is nonzero for only one of the two indices $\alpha=1,2$ (and $\rho(x,12)=\rho^{*}(x,21)=0$). In that case there is no condensate in this mean field approach.
III Variational approach: spatial invariance
@@
where $\hat{e}_{x},\hat{e}_{y}$ are unit vectors in the $x,y$ directions respectively, $h$ is a variational parameter, and $q=\pm 1$, where $q=+1$ induces an s-wave condensate, and $q=-1$ induces a d-wave condensate. We will denote the modified Hartree-Fock Hamiltonian as $H_{HFq}$. Let us regard the ground state of $H_{HFq}$, denoted $|\Omega_{h}\rangle$, as a trial wave functional, dependent on the variational parameter $h$. If the expectation value of energy density
$\mathcal{E}(h)={1\over L^{2}}\langle\Omega_{h}|H|\Omega_{h}\rangle$ (13)
of the full Hubbard Hamiltonian (1) is minimized at some ${h>0}$, then by the rules of the variational approach the ground state $|\Omega_{h}\rangle$ is a better approximation, as compared to the Hartree-Fock ground state $|\Omega_{0}\rangle$, to the true ground state of the Hubbard Hamiltonian. Then the procedure is to choose some set of parameters $\{U,\mu\}$ (setting $t=1$ for numerical work), and at each $U,\mu$ and variational parameter $h$ we compare $\mathcal{E}(0)$ to (i) $\mathcal{E}(h)$ for the ground state $|\Omega_{h}\rangle$ of $H_{HF-}$, associated with a d-wave condensate; and (ii) $\mathcal{E}(h)$ for the ground state $|\Omega_{h}\rangle$ of $H_{HF+}$, associated with an s-wave condensate. If $h>0$ is energetically favorable, in cases (i) and/or (ii), we choose the value of $h$ and sign of $q$ which gives the lowest $\mathcal{E}$. Whichever ground state has the lowest $\mathcal{E}$ is the state which is favored at the given $U,\mu$, at the electron density
$f={1\over L^{2}}\sum_{x}(\rho(x,11)+1-\rho(x,22))\ .$ (14)
We emphasize that the $h$ parameter is absent in the Hubbard Hamiltonian $H$, and only concerns the choice of trial ground state $|\Omega_{h}\rangle$.
We will say that the system is spatially homogeneous if the $\rho(x,\alpha\beta)=\rho(\alpha\beta)$ are position independent, which, within the Hartree-Fock approximation, is known to be true in lower density regions of the phase diagram. In this case the Hamiltonian can be diagonalized in momentum space, and the calculation is greatly simplified. In the absence of ferromagnetism there are equal numbers of spin up ($N_{\uparrow}$) and spin down ($N_{\downarrow}$) electrons, and therefore we have, in Nambu formalism,
$\rho(11)=1-\rho(22)\ ,$ (15)
with electron density
$f=\rho(11)+1-\rho(22)=2\rho(11)\ .$ (16)
It is also consistent to assume $\rho(12)=\rho(21)=0$. Translation invariance of $\rho$ implies that $[H]$ is diagonalized by momentum eigenstates, and on an $L\times L$ lattice, in the basis
@@
the matrix $[H]$ is decomposed, at each $k$ and for $q=-1$, into blocks
$[H]_{k}$ $=$ $\left[\begin{array}[]{cc}\epsilon^{1}_{k}&2h(\cos(k_{x})-\cos(k_{y}))\cr 2h(\cos(k_{x})-\cos(k_{y}))&\epsilon^{2}_{k}\end{array}\right]$ (21)
$+(U\rho(11)\rho(22)-\mu)\mathbbm{1}\ ,$
@@
The matrix $[H_{k}]$ is readily diagonalized. Denoting
$\epsilon_{k}$ $=$ $\epsilon^{1}_{k}=-2t((\cos(k_{x})+\cos(k_{y}))+Un_{\downarrow}-\mu$
$\beta_{k}$ $=$ $2h(\cos(k_{x})-q\cos(k_{y}))\ ,$ (23)
the eigenvalues are
$\lambda_{k\pm}=\pm\sqrt{\epsilon_{k}^{2}+\beta_{k}^{2}}+U\rho(11)\rho(22)-\mu\ ,$ (24)
and the eigenstate of lower energy is
$\varphi(k)$ $=$ $\left[\begin{array}[]{c}\varphi(k,1)\cr\varphi(k,2)\end{array}\right]$ (27)
$=$ ${\left[\begin{array}[]{c}\epsilon_{k}-\sqrt{\epsilon_{k}^{2}+\beta_{k}^{2}}\cr\beta_{k}\end{array}\right]\over\sqrt{\beta_{k}^{2}+\left(-\epsilon_{k}+\sqrt{\epsilon_{k}^{2}+\beta_{k}^{2}}\right)^{2}}}\ .$ (30)
The condensate (11) is
$P(k)=\varphi(k,1)\varphi(k,2)\ .$ (31)
As $h$ and $\beta_{k}\rightarrow 0$ the eigenvectors at $\epsilon_{k}\neq 0$ are simply
$\varphi(k)$ $=$ $\left[\begin{array}[]{c}0\cr 1\end{array}\right],~{}~{}~{}\epsilon_{k}>0$ (34)
$=$ $\left[\begin{array}[]{c}1\cr 0\end{array}\right],~{}~{}~{}\epsilon_{k}<0\ ,$ (37)
while at $\epsilon_{k}=0$ we have
$\varphi(k)$ $=$ ${1\over\sqrt{2}}\left[\begin{array}[]{c}-1\cr 1\end{array}\right],~{}~{}~{}\epsilon_{k}=0,~{}\beta_{k}/h>0$ (40)
$=$ ${1\over\sqrt{2}}\left[\begin{array}[]{c}1\cr 1\end{array}\right],~{}~{}~{}\epsilon_{k}=0,~{}\beta_{k}/h<0\ ,$ (43)
and in this limit
@@
$P(\mathbf{k})=\left\{\begin{array}[]{cl}\frac{1}{2}\text{sign}(\cos k_{x}-\cos k_{y})&\mathbf{k}\in\text{Fermi surface}\cr 0&\text{otherwise}\end{array}\right.\ .$ (46)
@@
We therefore seem to have a condensate even in the $h\rightarrow 0$ limit. This is due to an energy degeneracy on the Fermi surface ($\epsilon_{k}=0$) which is lifted by even an infinitesimal perturbation. Since this seems to be a case where the slightest perturbation breaks a ground state degeneracy, even as the strength of the breaking term goes to zero, it is tempting to regard this situation as an example of spontaneous symmetry breaking. In fact this view was already presented in Matsuyama:2022dtd , but there are two objections, both connected with lattice volume $L^{2}$. The first is that, unlike spontaneous symmetry breaking, this breaking exists already at finite volume; no special order of limits $L^{2}\rightarrow\infty$ followed by $h\rightarrow 0$ is necessary. But the more serious issue is that in the $h\rightarrow 0$ limit the condensate only exists exactly at the Fermi surface. In that case the condensate density is proportional to $1/L$, and vanishes in the infinite volume limit.
But there is no reason to take the $h\rightarrow 0$ limit. The real objective is to find the best possible approximation, within the limitations dictated by a single Slater determinant, to the ground state of the full Hubbard Hamiltonian of (1). The question is whether eigenstates of the modified Hartree-Fock Hamiltonian $H_{HFq}$ would lower the expectation value of the Hubbard Hamiltonian $H$ at some $h>0$. If so, then since one-particle wavefunctions would have non-zero amplitudes for both $\alpha$ index values, a condensate is obtained. At finite $h$ the condensate density does not vanish in the infinite volume limit, but is still concentrated in a small region of momentum space in the neighborhood of the Fermi surface. A typical result on a $16^{2}$ lattice is shown in Fig. 1.
Figure 1: Momentum space condensate $P(k)$ at $U=2,\mu=0.1$ and $h=0.002$, where we find density $f=0.72$. In this case, however, a lower energy is obtained at $h=0$.
But the question is whether the energy expectation value of the Hubbard Hamiltonian $\mathcal{E}(h)$ is less than $\mathcal{E}(0)$, for some $h\neq 0$, where the energy densities are
$\mathcal{E}(h)$ $=$ ${1\over L^{2}}\langle\Omega_{h}|H|\Omega_{h}\rangle$ (49)
$=$ $\sum_{k}\varphi^{\dagger}(k)\left[\begin{array}[]{cc}\epsilon^{1}_{k}&0\cr 0&\epsilon^{2}_{k}\end{array}\right]\varphi(k)+U\rho(11)\rho(22)-\mu\ ,$
and the $\varphi(k)$ are given in (30). Unfortunately this is not the case. No matter the choice of $U,\mu,h$, we always find ${\mathcal{E}(0)<\mathcal{E}(h)}$. Condensate states in the spatially homogenous region ($\rho(x,\alpha\beta)$ $x$-independent) are not energetically preferred, at least in this extension of the Hartree-Fock procedure.
IV Variational approach: spatial inhomogeneity
Although it is true that regions of low density in the $U$-density plane are spatially homogenous in the sense defined above, it is also well known that this is not generally the case at all densities, where stripes and other types of spatial inhomogenities in $\rho(x,\alpha\beta)$ are manifest. So the question is whether such inhomogeneities can change the disappointing conclusion of the previous section. To find out, we drop the assumption of spatial homogeneity and determine the ground state $|\Omega_{h}\rangle$ of the modified Hartree-Fock Hamiltonian $H_{HFq}$ self-consistently, by standard numerical methods. [^1] A useful check is to compare the results obtained by the methods of the previous section with the numerical approach that makes no assumption of spatially homogeneity, for values of $U,\mu$ where the latter method finds $\rho(x,\alpha\beta)$ to be constant. Since the answers must (and in fact do) agree in that case, this provides a modest check of our numerical code. Now
$|\Omega_{h}\rangle=\prod_{i=1}^{M}\bigg{(}\sum_{x_{i},\alpha_{i}}\phi^{h}_{i}(x_{i},\alpha_{i})\psi^{\dagger}_{\alpha_{i}}(x_{i})\bigg{)}|0\rangle\ .$ (51)
where
$\left([H]_{x\alpha,y\beta}+[V^{cond}]_{x\alpha,y\beta}\right)\phi^{h}_{i}(y,\beta)=E^{h}_{i}\phi^{h}_{i}(x,\alpha)\ .$ (52)
with
$\rho(x,\alpha\beta)$ $=$ $\langle\Omega_{h}|\psi^{\dagger}_{\alpha}(x)\psi_{\beta}(x)|\Omega_{h}\rangle$ (53)
$=$ $\sum_{i}\phi^{h*}_{i}(x,\alpha)\phi^{h}_{i}(x,\beta)\ ,$
determined self-consistently. This defines the trial state.
Defining also
$\rho_{nn}(x,\alpha\beta)$ $=$ $\sum_{i=1}^{L^{2}}\phi^{*h}_{i}(x,\alpha)\bigg{(}\phi^{h}_{i}(x+\hat{e}_{x},\beta)+\phi^{h}_{i}(x-\hat{e}_{x},\beta)$ (54)
$+\phi^{h}_{i}(x+\hat{e}_{y},\beta)+\phi^{h}_{i}(x-\hat{e}_{y},\beta)\bigg{)}\ ,$
we have
$\langle\Omega_{h}|H|\Omega_{h}\rangle=$
$~{}~{}~{}-t\sum_{x}\{\rho_{nn}(x,11)-\rho_{nn}(x,22)\}+U\sum_{x}\rho(x,11)$
$~{}~{}~{}-U\sum_{x}\bigg{(}\rho(x,11)\rho(x,22)-\rho(x,12)\rho(x,21)\bigg{)}$
$~{}~{}~{}-\mu\sum_{x}\{\rho(x,11)+1-\rho(x,22)\}\ ,$ (55)
and energy density
$\mathcal{E}(h)={1\over L^{2}}\langle\Omega_{h}|H|\Omega_{h}\rangle\ .$ (56)